if 48 grams of Al and 100 g of Br2 are reacted, how much AlBr3 can be produced
1
Expert's answer
2020-06-18T14:41:42-0400

2Al+3Br2=2AlBr3

1.we can find out which substance is left over.for this we first find the amount of Al and Br2 in (mol)

n(Al)=48/27=1.778mol

n(Br2)=100/160=0.625mol

2.according to the above calculations, Al increases. Therefore, we can find AlBr3 based on the amount (mol) of Br2:

3mol Br2——> 2 mol AlBr3

0.625mol Br2——> x mol AlBr3

x=0.625*2/3=0.4167 mol AlBr3

3.now we find the mass of AlBr3. for this we use the following formula: m = n * M

m=0.4167*267=111.25 gr AlBr3

Answer: 111.25 gr